Circular No. 6554

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1997 B1 (KOBAYASHI) Elliptical orbital elements from the positions on IAUC 6553:

T = 1997 Mar. 4.295 TT  ω = 184.^{o}202
e = 0.76831  Ω = 329.032 }2000.0
q = 2.06219 AU  i = 12.490

a = 8.90047 AU n^{o} = 0.037118 P = 26.55 years

1997 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{2}
Feb. 1 9^{h}59.^{m}69 +16^{o}20.^{′}4 1.120 2.088 165.^{o}4 6.^{o}8 17.4
6 9 56.88 +15 42.4 1.101 2.081 170.9  4.3 17.2
11 9 53.77 +15 03.3 1.088 2.074 176.2  1.8 17.0
16 9 50.54 +14 22.9 1.082 2.069 177.2  1.3 16.9
21 9 47.36 +13 41.3 1.082 2.066 172.0  3.8 17.1
26 9 44.40 +12 58.7 1.088 2.063 166.5  6.4 17.2
Mar. 3 9 41.83 +12 15.3 1.099 2.062 161.1  9.0 17.4
8 9 39.76 +11 31.3 1.117 2.063 155.8 11.4 17.5
13 9 38.33 +10 46.8 1.140 2.064 150.6 13.7 17.6

S. Nakano, Sumoto, Japan, reports that images obtained at Kuma Kogen by A. Nakamura (0.6-m f/6 Ritchey-Chrétien) on Feb. 4.75 UT show an obvious coma and a tail 35′′ long in p.a. 290^{o}.

SUPERNOVA 1997X IN NGC 4691 S. Benetti, M. Turatto and I. Perez, European Southern Observatory;

and L. Wisotzki, Hamburger Sternwarte, communicate: “Inspection of a CCD spectrum (range 316–1000 nm, resolution 2.0 nm), obtained on Feb. 4.39 UT with the 1.52-m telescope (+ Boller & Chivens spectrograph) at La Silla, confirms that supernova 1997X is of type Ic and about one week after maximum. The spectrum, which shows a quite red continuum, is dominated by lines (with P-Cyg profile) of Ca II (H and K + infrared triplet) and Fe II. A strong line (also with P-Cyg profile, maximum measured at 585.0 nm, minimum at 566.3 nm) could be identified with a blend of He I 587.6-nm and Na I D. Other He I lines could be present at (mimima) 646.0 nm, 685.5 and possibly 479.5 nm. If the helium identification is confirmed (the helium lines becoming more intense at later phases), the expansion velocity deduced for the helium-emitting layer, ∼ 10 000 km/s, is smaller than in the type Ic SNe 1987M and 1994I (Clocchiatti et al. 1996, Ap.J. 462, 462). The Ca II expansion velocity is ∼ 16 000 km/s. The O I 777.4-nm line is also present.”

1997 February 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden